### v1

@@ 5.3 Symmetric Randomized Hadamard Transform
where applying ${\mathbf{Q}}$ on both sides of ${\mathbf{A}}$ is crucial to maintaining the PSD structure of the system. Recall that the transform can be defined as ${\mathbf{Q}}={\mathbf{H}}{\mathbf{D}}$, where ${\mathbf{D}}=\frac{1}{\sqrt{n}}\mathrm{diag}(d_{1},...,d_{n})$ and $d_{i}$ are independent random ±$\pm 1$ signs (Rademacher variables). So, the dominant cost of this preprocessing step is applying the Hadamard transform ${\mathbf{H}}$ on both sides of the matrix ${\mathbf{D}}{\mathbf{A}}{\mathbf{D}}$. The classical recursive algorithm for doing this, which we refer to as the Fast Hadamard Transform (FHT, see Appendix C), takes $mn\log n$ operations to compute ${\mathrm{FHT}}({\mathbf{M}})={\mathbf{H}}{\mathbf{M}}$ for an $n\times m$ matrix ${\mathbf{M}}$. Thus, the cost of computing ${\mathbf{Q}}{\mathbf{A}}{\mathbf{Q}}^{\scriptscriptstyle{\top}}={\mathrm{FHT}}({\mathrm{FHT}}({\mathbf{D}}{\mathbf{A}}{\mathbf{D}})^{\scriptscriptstyle{\top}})$ is roughly $2n^{2}\log n$ operations. This suggests that, in order to maintain the positive definite structure for CD++, we must double the preprocessing cost compared to Kaczmarz++.
We show that this trade-off can be entirely avoided: By exploiting the symmetric structure of ${\mathbf{A}}$, we perform the two FHTs simultaneously at the cost of one FHT applied to a general matrix. We achieve this using a specialized recursive algorithm, which we call SymFHT (Algorithm 2). Here, for simplicity we write the recursion assuming that the input matrix is at least $2\times 2$. Naturally, the base case of the recursion is a $1\times 1$ input matrix, in which case we let ${\mathrm{SymFHT}}({\mathbf{A}})={\mathbf{A}}$.
@@
1:function SymFHT(${\mathbf{A}}$) $\triangleright$ Input: Symmetric matrix ${\mathbf{A}}=\begin{bmatrix}{\mathbf{A}}_{11}&{\mathbf{A}}_{12}\\ {\mathbf{A}}_{12}^{\scriptscriptstyle{\top}}&{\mathbf{A}}_{22}\end{bmatrix}$.
@@
Note that, given an $n\times n$ matrix ${\mathbf{A}}$ broken down into four $n/2\times n/2$ blocks, the function SymFHT performs two recursive calls corresponding to the two diagonal blocks ${\mathbf{A}}_{11}$ and ${\mathbf{A}}_{22}$, since both of these blocks are symmetric. The two off-diagonal blocks ${\mathbf{A}}_{12}$ and ${\mathbf{A}}_{12}^{\scriptscriptstyle{\top}}$ are not symmetric, so we must revert back to applying the classical FHT twice. But crucially, since the off-diagonal blocks are identical up to a transpose, we only have to transform one of them, which gives our recursion its computational gain, as shown in the following result. See Appendix C for proof.
@@
Given an $n\times n$ symmetric matrix ${\mathbf{A}}$, where $n$ is a power of $2$, Algorithm 2 returns ${\mathbf{H}}{\mathbf{A}}{\mathbf{H}}$ after at most $n^{2}(2.5+\log n)$ arithmetic operations.
@@ 5.4 Error Estimation and Adaptive Tuning
Combining the above ideas, we obtain CD++, given in Algorithm 3.
@@
3:${\mathbf{A}}\leftarrow{\mathrm{SymFHT}}({\mathbf{D}}{\mathbf{A}}{\mathbf{D}})$, ${\mathbf{b}}\leftarrow{\mathrm{FHT}}({\mathbf{D}}{\mathbf{b}})$; $\triangleright$ See Algorithm 2.
4:Initialize ${\mathbf{m}}_{0}\leftarrow\mathbf{0}$, $\rho\leftarrow 0$, $\eta\leftarrow\frac{s}{2n}$, ${\mathcal{B}}\leftarrow\emptyset$, $\tau\leftarrow\lceil n/s\rceil$, ${\mathcal{E}}_{0}\leftarrow{\mathcal{E}}_{1}\leftarrow 0$, $r\leftarrow 0$;
@@
14:     if $t<\tau\mod 2\tau$ then ${\mathcal{E}}_{0}\leftarrow{\mathcal{E}}_{0}+\|{\mathbf{r}}_{t}\|^{2}$; else ${\mathcal{E}}_{1}\leftarrow{\mathcal{E}}_{1}+\|{\mathbf{r}}_{t}\|^{2}$;
15:     if $t=2\tau-1\mod 2\tau$ then
@@
18:         $\rho\leftarrow 1-r^{1/\tau}$; $\triangleright$ Convergence rate estimate.
@@ 6 Numerical Experiments
In this section, we present numerical experiments that support our theory. First, we demonstrate that the convergence analysis carried out in Sections 2-4 accurately predicts the performance of Kaczmarz++/CD++, by evaluating the effect of adaptive acceleration, block memoization and regularized projections on their convergence rate. Then, we compare our methods to popular Krylov solvers, CG [HS52] and GMRES [SS86], showing that the Kaczmarz++ framework achieves better computational cost for certain classes of linear systems that arise naturally in applications such as machine learning, as suggested by our theory.
@@ 6.1 Experimental Setup
We set up our experiments to evaluate how well the solvers exploit large outlying singular values to achieve fast convergence for ill-conditioned systems. To that end, we consider two families of positive semidefinite linear systems which naturally exhibit such spectral structure, resulting in 20 different linear system tasks (see Appendix D for formal definitions).
@@
To gain precise control on the eigenvalue distribution of the system, we first consider a collection of synthetic benchmark matrices with a bell-shaped spectrum, constructed via the make_low_rank_matrix function in Scikit-learn [PVG+11]. We control the number of large outlying eigenvalues via the function parameter effective_rank, choosing among four values (25, 50, 100, and 200).
@@
We also consider four real-world benchmark datasets, available through Scikit-learn [PVG+11] and OpenML [VvRBT13]: Abalone, California_housing, Covtype, and Phoneme. We transform these datasets using a kernel function to produce a PSD kernel matrix. For the kernel function, we consider two types of radial basis functions (Gaussian and Laplacian), each with two different values of width $\gamma\in\{0.1,0.01\}$. These are known to produce highly ill-conditioned matrices with fast spectral decays, leading to many large outlying eigenvalues [RW06].
Each resulting PSD matrix ${\mathbf{A}}$ is truncated to dimensions $4096\times 4096$. Then, following standard applications in Kernel Ridge Regression [AM15], for each matrix we define a regularized linear system $({\mathbf{A}}+\phi{\mathbf{I}}){\mathbf{x}}={\mathbf{b}}$, where ${\mathbf{b}}$ is a standard Gaussian vector and $\phi=0.001$. We evaluate the returned solutions via the normalized residual:
$\epsilon=\|({\mathbf{A}}+\phi{\mathbf{I}}){\mathbf{x}}-{\mathbf{b}}\|/\|{\mathbf{b}}\|.$ (6.1)
@@ 6.2 Verifying Our Convergence Analysis
To verify our convergence analysis, we implement four variants of our algorithms and plot the per-iteration convergence on all tasks (details in Appendix D).
Solver Accelerate Memoize
CD
CD+Accel
CD+Memo
Full CD++
Table 1: Implemented variants of CD++.
Since the tasks are all positive semidefinite, we focus on CD++  (Algorithm 3). To isolate the effect of its individual components, such as block size, acceleration, and memoization, we consider four variants of the method that toggle acceleration and memoization on/off, as illustrated in Table 1, each with four different block sizes (25, 50, 100, and 200). In particular, no memoization means that we sample a new block set $S$ uniformly at random from (${[n]\choose s}$ at every step (and compute its Cholesky factor), while no acceleration means setting the momentum step size $\eta$ to 00. We show convergence plots for synthetic matrices with effective rank $100$ and $200$ in Figure 1, and include the remaining 6 plots (2 for synthetic matrices and 4 for the Abalone dataset) in Figure 3 of Appendix D.
Figure 1: Convergence plots for different variants of CD++ (see Table 1) using four block sizes (25, 50, 100, and 200), on synthetic test matrices with effective rank 100 (top) and 200 (bottom). Note that the iterations range in each column is scaled to keep (iterations×$\times$×block-size) consistent.
@@
First, observe that as we increase the block size, all of the methods achieve faster per-iteration convergence rate, as expected from our theory. Looking closely, we can see that this rate improves more than just proportionally to the increase in block size, which means that larger blocks lead to greater efficiency in terms of how many rows of ${\mathbf{A}}$ need to be processed by the algorithm to reach certain accuracy. This corresponds to the $\bar{\kappa}_{k}$ condition number in our theory, which decreases as the block size increases, indicating that the methods do exploit large outlying eigenvalues. Also, as we increase the effective rank of the matrix (i.e., the number of large eigenvalues, corresponds to $k$ in our theory), we need larger block sizes to attain fast convergence.
@@
Comparing CD and CD+Accel, we see that our adaptive acceleration significantly improves the convergence rate, particularly for small-to-moderate block sizes. However, when the block size is much larger than the effective rank, then the benefit of acceleration will necessarily be reduced (top right plot), because in this case the condition number $\bar{\kappa}_{k}$ is just a small constant.
@@
Comparing CD+Accel and Full CD++, we see that introducing our block memoization scheme (i.e., sampling from a small collection of random blocks) only slightly reduces the per-iteration convergence rate compared to the method that samples a new random block at every step. This is explained by our analysis in Theorem 4.2. The benefits of block memoization become apparent once we compare the computational cost of both procedures, as discussed below (Figure 2).
@@
In Appendix D.3, we tested Full CD++ with different choices of the projection regularizer $\lambda$. In all our experiments, the convergence remained largely unchanged for any $\lambda\in[0,0.01]$ (note that our theory requires $\lambda>0$). This suggests that regularizing the Kaczmarz projection can be done without sacrificing the convergence, and so, we recommend using CD++ with a small but positive $\lambda$ to ensure stable computation of the Cholesky factors (we used $\lambda=10^{-8}$ as a default).
@@
Figure 2: Computational cost comparison, measuring floating point operations (FLOPs) needed to reach a given error threshold on four kernel matrices constructed from the Abalone dataset. Above, we show total FLOPs for GMRES and CD++ to reach one of two error thresholds, $\epsilon\in\{10^{-4},10^{-8}\}$. Below, we show convergence-vs-FLOPs plots, including CG and CD+Accel as additional baselines.
@@ 6.3 Comparison with Krylov Subspace Methods
Next, we evaluate the computational cost of our methods, comparing them to Krylov solvers. Specifically, we count floating point operations (FLOPs) needed to reach a given error threshold. Here, we show the results for kernel matrices obtained from the Abalone dataset, while results for the remaining test matrices are in Appendix D.2.
In Figure 2 (bottom) we show the convergence plots of the methods, with FLOPs instead of iterations on the x-axis (for our methods, this includes the cost of RHT pre-processing). We again compare Full CD++ and CD+Accel (i.e., Algorithm 3 with and without block memoization, using block size 200). We see that, even though block memoization slightly reduces the per-iteration convergence (Figure 1), it more than makes up for this in the computations. In particular, Full CD++ exhibits a phase transition where it accelerates past CD+Accel once enough blocks have been memoized and it no longer pays for the block Cholesky factorizations (e.g., see bottom left plot in Figure 2).
We compare the computational cost of CD++ with two Krylov solvers, CG and GMRES. First, we observe that CG struggles to converge on all of the tested matrices. This suggests that the systems are indeed ill-conditioned, and CG cannot overcome this effectively due to numerical stability issues. On the other hand, GMRES avoids these issues by maintaining an explicit Krylov basis, and thus after a number of initial iterations, it exploits the large outlying eigenvalues to achieve fast convergence.
@@
These overall trends are reflected in Figure 2 (top), showing the total FLOP counts of GMRES and CD++ with two error thresholds, $\epsilon\in\{10^{-4},10^{-8}\}$. We see that for $\epsilon=10^{-4}$, CD++ is consistently more efficient than GMRES, while for $\epsilon=10^{-8}$, GMRES overtakes CD++ in some cases. Across all tested matrices (see Appendix D.2 for details), we observed that for $\epsilon=10^{-4}$, CD++ performed better than GMRES in 18 out of 20 cases, while for $\epsilon=10^{-8}$, it did so in 14 out of 20 cases.

### v2

@@ 5.3 Symmetric Randomized Hadamard Transform
where applying ${\mathbf{Q}}$ on both sides of ${\mathbf{A}}$ is crucial to maintaining the PSD structure of the system. Recall that the transform can be defined as ${\mathbf{Q}}={\mathbf{H}}{\mathbf{D}}$, where ${\mathbf{D}}=\frac{1}{\sqrt{n}}\mathrm{diag}(d_{1},...,d_{n})$ and $d_{i}$ are independent random ±1$\pm$ 1± 1 signs (Rademacher variables). So, the dominant cost of this preprocessing step is applying the Hadamard transform ${\mathbf{H}}$ on both sides of the matrix ${\mathbf{D}}{\mathbf{A}}{\mathbf{D}}$. The classical recursive algorithm for doing this, which we refer to as the Fast Hadamard Transform (FHT, see Appendix E), takes $mn\log n$ operations to compute ${\mathrm{FHT}}({\mathbf{M}})={\mathbf{H}}{\mathbf{M}}$ for an $n\times m$ matrix ${\mathbf{M}}$. Thus, the cost of computing ${\mathbf{Q}}{\mathbf{A}}{\mathbf{Q}}^{\scriptscriptstyle{\top}}={\mathrm{FHT}}({\mathrm{FHT}}({\mathbf{D}}{\mathbf{A}}{\mathbf{D}})^{\scriptscriptstyle{\top}})$ is roughly $2n^{2}\log n$ operations. This suggests that, in order to maintain the positive definite structure for CD++, we must double the preprocessing cost compared to Kaczmarz++.
We show that this trade-off can be entirely avoided: By exploiting the symmetric structure of ${\mathbf{A}}$, we perform the two FHTs simultaneously at the cost of one FHT applied to a general matrix. We achieve this using a specialized recursive algorithm, which we call SymFHT (Algorithm 2). Here, for simplicity we write the recursion assuming that the input matrix is at least $\times 2$ 2. Naturally, the base case of the recursion is a $\times 1$ 1 input matrix, in which case we let ${\mathrm{SymFHT}}({\mathbf{A}})={\mathbf{A}}$.
@@
[…]
@@
Note that, given an $n\times n$ matrix ${\mathbf{A}}$ broken down into four $n/2\times n/2$ blocks, the function SymFHT performs two recursive calls corresponding to the two diagonal blocks ${\mathbf{A}}_{11}$ and ${\mathbf{A}}_{22}$, since both of these blocks are symmetric. The two off-diagonal blocks ${\mathbf{A}}_{12}$ and ${\mathbf{A}}_{12}^{\scriptscriptstyle{\top}}$ are not symmetric, so we must revert back to applying the classical FHT twice. But crucially, since the off-diagonal blocks are identical up to a transpose, we only have to transform one of them, which gives our recursion its computational gain, as shown in the following result. See Appendix E for proof.
@@
Given an $n\times n$ symmetric matrix ${\mathbf{A}}$, where $n$ is a power of 222, Algorithm 2 returns ${\mathbf{H}}{\mathbf{A}}{\mathbf{H}}$ after at most $n^{2}(2.5+\log n)$ arithmetic operations.
@@ 5.4 Error Estimation and Adaptive Tuning
Combining the above ideas, we obtain CD++, given in Algorithm 3. Also, Algorithm 5 in the supplement provides the complete pseudocode for Kaczmarz++, including the above adaptive tuning scheme as well as block memoization with preconditioned LSQR from Section 4.
@@
[…]
4:Initialize ${\mathbf{m}}_{0}\leftarrow\mathbf{0}$, $\rho\leftarrow 0$, $\eta\leftarrow\frac{s}{2n}$, ${\mathcal{B}}\leftarrow\emptyset$, $\zeta\leftarrow\lceil n/s\rceil$, ${\mathcal{E}}_{0},{\mathcal{E}}_{1}\leftarrow 0$;
@@
14:  if $t<\zeta\mod 2\zeta$ then ${\mathcal{E}}_{0}\leftarrow{\mathcal{E}}_{0}+\|{\mathbf{r}}_{t}\|^{2}$; else ${\mathcal{E}}_{1}\leftarrow{\mathcal{E}}_{1}+\|{\mathbf{r}}_{t}\|^{2}$;
15:  if $t=2\zeta-1\mod 2\zeta$ then
@@
18:   $\rho\leftarrow 1-r^{1/\zeta}$; $\triangleright$ Convergence rate estimate.
@@ 6 Numerical Experiments
In this section, we present numerical experiments that support our theory. First, we demonstrate that the convergence analysis carried out in Sections 2-4 accurately predicts the performance of Kaczmarz++ on a collection of synthetic linear system tasks, by evaluating the effect of adaptive acceleration, block memoization and inexact projections on the convergence rate. Then, focusing on a real-world positive definite system task, we evaluate CD++ against popular Krylov solvers, CG [HS52] and GMRES [SS86], showing that our algorithmic framework achieves better computational cost for certain classes of linear systems that arise naturally in applications such as machine learning, as suggested by our theory. We note that due to the limitations in our computational setup, we evaluate our algorithms on moderately sized matrices, which showcases our theoretical contributions, and do not report running times or test against direct solvers.
@@ 6.1 Experimental Setup
We set up our experiments to evaluate how well the solvers exploit large outlying singular values to achieve fast convergence for ill-conditioned systems. To that end, we consider two families of linear systems which naturally exhibit such spectral structure (see Appendix F for formal definitions).
@@
To gain precise control on the eigenvalue distribution of the system, we first consider a collection of synthetic benchmark matrices with a bell-shaped spectrum, constructed via the make_low_rank_matrix function in Scikit-learn [PVG+11]. We control the number of large outlying singular values via the parameter effective_rank, choosing among four values (25, 50, 100 and 200).
@@
We also consider four real-world benchmark datasets, available through Scikit-learn [PVG+11] and OpenML [VvRBT13]: Abalone, California_housing, Covtype, and Phoneme. We transform these datasets using a kernel function to produce a PSD kernel matrix. For the kernel function, we consider two types of radial basis functions (Gaussian and Laplacian), each with two different values of width $\gamma\in\{0.1,0.01\}$. These are known to produce highly ill-conditioned matrices with fast spectral decays, leading to many large outlying eigenvalues [RW06]. Each resulting PSD matrix is truncated to dimensions $\times 4096$ 4096. Then, following standard applications in Kernel Ridge Regression [AM15], we augment the matrix with a regularization term $\phi{\mathbf{I}}$, where $\phi=0.001$.
For each resulting test matrix ${\mathbf{A}}$, we solve the linear system ${\mathbf{A}}{\mathbf{x}}={\mathbf{b}}$, where ${\mathbf{b}}$ is a standard Gaussian vector, and we evaluate the estimates via the normalized residual:
$\epsilon=\|{\mathbf{A}}{\mathbf{x}}-{\mathbf{b}}\|/\|{\mathbf{b}}\|.$ (6.1)
@@ 6.2 Verifying Our Convergence Analysis
To verify our convergence analysis, we implement four variants of Kaczmarz++ and plot the per-iteration convergence on ill-conditioned synthetic low-rank matrices (details in Appendix F). Note that for these experiments we are testing Kaczmarz++ on rectangular 4096×10244096$\times$ 10244096 × 1024 linear systems (constructed via the make_low_rank_matrix function in Scikit-learn [PVG+11]).
Solver Name Accelerate Memoize
Kaczmarz
K++ w/o Accel
K++ w/o Memo
Full K++
Table 1: Evaluated variants of Kaczmarz++. Analogous names apply in Section 6.3 with CD++ instead of K++.
Inner solver. First, we test the influence of computing the inner projection steps inexactly via sketch-and-precondition LSQR (Section 4.1). As shown in Figure 1, even very few steps of LSQR suffice to attain the fast linear convergence of Kaczmarz++. With only 888 steps of LSQR, there is barely any difference between Kaczmarz++ and the alternative using exact projections. This is supported by our theory: as discussed in Lemma 4.1, by using a preconditioner constructed via SRHT with sketch size proportional to the block size, the condition number of the arising sub-problem can be reduced to 2$\leq$ 2≤ 2. In our experiments, we observed that using sketch size $\tau$ equal twice the block size is sufficient.
Next, to isolate the effect of its individual components (such as block size, acceleration, and memoization), we consider four variants of Kaczmarz++ that toggle acceleration and memoization on/off, as illustrated in Table 1. In Figure 2, we show convergence plots for synthetic matrices with effective rank 505050 and 100100100, each with block sizes 100100100 and 200200200 (the plots for matrices with effective rank 25 and 200 are in Appendix F.1). In particular, no memoization means that we sample a new block set $S$ uniformly at random from (${[m]\choose s}$ at every step (and compute its Cholesky factor), while no acceleration means setting the momentum step size $\eta$ to 0.
Figure 1: Convergence plots for Kaczmarz++ varying the number of steps of preconditioned LSQR in each regularized projection step, with two block sizes (100, 200) on synthetic test matrices with effective rank 50 (left) and 100 (right). K++(LSQR-X) is Kaczmarz++ using X steps of LSQR, while K++(Cholesky) means Kaczmarz++ with exact inner solver using Cholesky decomposition.
Figure 2: Convergence plots for different variants of Kaczmarz++ (see Table 1) using two block sizes (100, 200), on synthetic test matrices with effective rank 50 (left) and 100 (right). Note that the iteration range in each plot is scaled to keep (iterations×$\times$×block-size) consistent.
@@
First, observe that as we increase the block size, all of the methods achieve faster per-iteration convergence rate, as expected from our theory. Looking closely, we can see that this rate improves more than just proportionally to the increase in block size, which means that larger blocks lead to greater efficiency in terms of how many rows of ${\mathbf{A}}$ need to be processed by the algorithm to reach certain accuracy. This corresponds to the $\bar{\kappa}_{k}$ condition number in our theory, which decreases as the block size increases, indicating that the methods do exploit large outlying singular values. Also, as we increase the effective rank of the matrix (i.e., the number of large singular values, $k$ in our theory), we need larger block sizes to attain fast convergence.
@@
Comparing Full K++ and K++ w/o Accel, we see that our adaptive acceleration significantly improves the convergence rate, particularly for the smaller block size 100. Note that if the block size is chosen to be much larger than the effective rank, then the benefit of acceleration will necessarily get diminished, because in this case the condition number $\bar{\kappa}_{k}$ is just a small constant.
@@
Comparing Full K++ and K++ w/o Memo, we see that introducing our block memoization scheme (i.e., sampling from a small collection of random blocks) only slightly reduces the per-iteration convergence rate compared to the method that samples a new random block at every step. This is explained by our analysis in Theorem 4.2. The benefits of block memoization become apparent once we compare the computational cost of both procedures, as discussed below (Figure 3).
@@
In Appendix F.3, we also tested the effect of different choices of the projection regularizer $\lambda$ on our methods. In all our experiments, the convergence remained largely unchanged for any $\lambda\in[0,0.01]$ (our theory requires $\lambda>0$). This suggests that regularizing the Kaczmarz projection can be done without sacrificing convergence, and so, we recommend using a small but positive $\lambda$ to ensure stable computation of the Cholesky factors (we used $\lambda=10^{-8}$ as a default).
@@
Figure 3: Computational cost comparison, measuring floating point operations (FLOPs) needed to reach a given error threshold on four kernel matrices constructed from the Abalone dataset. Above, we show total FLOPs for GMRES and CD++ to reach one of two error thresholds, $\epsilon\in\{10^{-4},10^{-8}\}$. Below, we show convergence-vs-FLOPs plots, including CG and CD+Accel as additional baselines. We also include CD++ w/o RHT to test the influence of RHT on the convergence.
@@ 6.3 Comparison with Krylov Subspace Methods
Next, we evaluate the computational cost of our methods, comparing them to Krylov solvers. Specifically, we count floating point operations (FLOPs) needed to reach a given error threshold for the positive definite linear systems arising from benchmark kernel matrices in machine learning. Here, we show the results for the Abalone dataset, while results for the remaining test matrices are in Appendix F.2. In Figure 3 (bottom) we show the convergence plots of the methods, with FLOPs instead of iterations on the x-axis (this includes the cost of RHT pre-processing, when appropriate). Since the tasks are positive definite, we focus on evaluating different variants of CD++.
First, we compare Full CD++ and CD++ w/o Memo (i.e., Algorithm 3 with and without block memoization, using block size 200). We see that, even though block memoization slightly reduces the per-iteration convergence (Figure 2), it more than makes up for this in the computations. In particular, Full CD++ exhibits a phase transition where it accelerates past CD++ w/o Memo once enough blocks have been memoized and it no longer pays for the block Cholesky factorizations.
We also evaluate the effect of RHT preprocessing on the overall computational cost of the method by comparing Full CD++ and CD++ w/o RHT. The preprocessing cost itself is relatively negligible, as can be seen by how much the beginning of the convergence curve of Full CD++ is shifted away from 0 (by around 0.2e9 FLOPS). Furthermore, RHT provides a substantial gain in overall performance for roughly half of the test matrices, and this can be seen consistently across the remaining plots in the supplement. Importantly, even without RHT preprocessing, CD++ exhibits good convergence for most of the problems in our testing pool. These results suggest the possibility that CD++ w/o RHT may also be effective for sparse systems, where RHT preprocessing is not desirable. However, further empirical evidence on large-scale sparse systems could be interesting future work to address this possibility.
Finally, we compare the computational cost of CD++ with two Krylov solvers [^2] We use the SciPy implementation of CG [VGO+20] and the PyAMG implementation of GMRES [BOSS23], which were chosen based on how amenable they are to our implementation of FLOPs counting., CG and GMRES. First, we observe that CG struggles to converge on all of the tested matrices. This suggests that the systems are indeed ill-conditioned, and CG cannot overcome this effectively due to numerical stability issues. On the other hand, GMRES avoids these issues by maintaining an explicit Krylov basis, and thus after a number of initial iterations, it exploits the large outlying eigenvalues to achieve fast convergence.
@@
These overall trends are reflected in Figure 3 (top), showing the total FLOP counts of GMRES and CD++ with two error thresholds, $\epsilon\in\{10^{-4},10^{-8}\}$. We see that for $\epsilon=10^{-4}$, CD++ is consistently more efficient than GMRES, while for $\epsilon=10^{-8}$, GMRES overtakes CD++ in some cases. Across all tested matrices (see Appendix F.2 for details), we observed that for $\epsilon=10^{-4}$, CD++ performed better than GMRES in 18 out of 20 cases, while for $\epsilon=10^{-8}$, it did so in 14 out of 20 cases. Interestingly, CD++ w/o RHT outperformed GMRES on all 20 test matrices for $\epsilon=10^{-4}$, while for $\epsilon=10^{-8}$, it did so in 9 out of 20 cases. This suggests that RHT preprocessing provides a significant computational benefit mainly in the high-precision regime, and that it might be preferable to skip this step when moderate precision is sufficient.